### v1

# Exploring Next Token Prediction For Optimizing Databases

Abstract.
The Next Token Prediction paradigm (NTP, for short) lies at the forefront of modern large foundational models that are pre-trained on diverse and large datasets. These models generalize effectively and have proven to be very successful in Natural Language Processing (NLP). Inspired by the generalization capabilities of Large Language Models (LLMs), we investigate whether the same NTP paradigm can also be applied to DBMS design and optimization tasks. Adopting NTP directly for database optimization is non-trivial due to the fundamental differences between the domains. In this paper, we present a framework termed Probe and Learn (PoLe) for applying NTP to optimize database systems. PoLe leverages Decision Transformers and hardware-generated tokens to effectively incorporate NTP into database systems. Preliminary results from the main-memory index scheduling task demonstrate that adopting NTP can improve both performance and generalizability.
Next Token Prediction; Database Optimization
1. Introduction
Hardware and workload are the two dominant factors that have shaped the DBMS research over the past 50 years, and are termed as the “Game Changers” in databases (Ailamaki, 2021). Both factors are rapidly evolving and consistently reshaping the design principles and optimization techniques in modern DBMSs. On the hardware front, recent examples include works on GPUs, e.g., (Boeschen and Binnig, 2022; Mageirakos et al., 2022; Sioulas et al., 2019; Chrysogelos et al., 2019; He and Yu, 2011; Breß, 2013), Persistent Memory (PMEM), e.g., (Zuo et al., 2018; van Renen et al., 2018), SSDs, e.g., (Tsirogiannis et al., 2009; Shah et al., 2008), and Compute Express Link (CXL), e.g., (Ahn et al., 2022; Lee et al., 2024). Recently, ARM processors have made significant progress in high performance computing, and are becoming competitive with the x86 processors with high-end servers, e.g., Apple Silicon (Wikipedia, 2024), Amazon Graviton (Amazon, 2024b), and NVIDIA Grace (NVIDIA, 2024). On the workload front, the landscape is equally dynamic, and has taken a drastic turn over the recent years, with the rise of Large Language Model (LLM, for short) workloads. LLM workloads impose distinct computational and memory demands, further pushing the boundaries of DBMS design and optimization (Liu et al., 2024).
Another game-changing factor is the heterogeneity of both the hardware and the workloads. On the hardware front, this heterogeneity manifests in diverse compute and memory chips. For example, modern servers are equipped with large cores, small cores, or a mixture of both. Larger cores are preferable for serialized execution, whereas parallelization suits smaller cores. Likewise, the same Integrated Memory Controllers (IMC) now-a-days support both DRAM and PMEM. In addition, the heterogeneity leads to variable read and write access latencies across hardware, e.g., in Non Uniform Memory Access (NUMA) servers accessing data that resides in a remote DRAM chip can incur up to 3×3$\times$3 × more latency than accessing data that resides in the local DRAM chip (Li et al., 2013). On the workload front, this heterogeneity stems from the diverse range of applications that DBMSs must support, including OLTP, OLAP, HTAP, and modern ML workloads. Each of these hardware and workload applications has distinct performance characteristics, demanding tailored optimization strategies. Adds to this is the resurgence of cloud databases taking over the traditional on-site DBMSs. Amazon Elastic Compute Cloud (EC2) (Amazon, 2024a) provides as many as 750 instance choices equipped with a diverse range of processors, memory, storage, and networks hardware to fit different workload applications.
To address these challenges, DBMS engines need to generalize across heterogeneous hardware and workload applications without sacrificing performance. We need a complete paradigm shift in how DBMSs are designed, one that embraces hardware and workload heterogeneity to gain best performance, while being able to generalize and adapt across new hardware architectures and workload applications without manual intervention. Rather than constantly reinventing the wheel to keep pace with the ever-changing hardware and workload landscape, the goal is to distill decades of DBMS knowledge for diverse workloads and hardware into a compact, transferable form that is both generalizable and adaptable.
Large Language Models (LLMs) have proven to be very successful in generalization tasks across multiple fields, including but not limited to Natural Language Processing (NLP), Computer Vision, and Robotics. The Next Token Prediction (NTP) lies at the core of LLMs. It is a foundational concept in modern AI that involves training models to predict the next element in a sequence based on the preceding context. By mastering NTP, LLMs are responsible for a major share of all AI breakthroughs in tasks such as machine translation, text summarization, code generation, over the past decade. It is only natural to question how this simple yet powerful NTP paradigm can be applied to DBMS design and optimizations, and whether the NTP Paradigm can help in building data systems that are capable of generalizing across diverse hardware and workload applications while maintaining high performance.
In this paper, we examine the feasibility of applying the Next Token Prediction paradigm in the context of database systems (cf. § 2). We address the challenges associated with adopting this paradigm for database systems, and identify two building blocks, i.e., Decision Transformers (DT, for short) and DB-tokens, i.e., hardware profiles that are essential in translating NTP to DBMS optimizations (cf. § 3). We propose a general framework, Probe and Learn (PoLe, for short) for database systems to effectively incorporate the NTP paradigm (cf. § 4), and evaluate its effectiveness through a case-study on the main-memory index scheduling task. Preliminary results on the main-memory index scheduling demonstrate that the PoLe framework can learn novel scheduling techniques improving the index performance in query throughput up to 3$\times$ (cf. § 5).
2. Database Systems And The Next Token Prediction Paradigm
Next Token Prediction: At a high level, the next token prediction (NTP, for short) models the probability of the next token in a given sequence based on the past tokens. In Natural Language Processing (NLP), each token is a discrete unit, i.e., a word, a sub-word, or a character. For any sequence $\bm{\tau}$, let $\tau_{i}$ denote the $i$-th token, and let $\bm{\tau_{<i}}$ denote the $i-1$ tokens preceding $\tau_{i}$. Then, the goal of NTP is to estimate the probability distribution of $\tau_{i}$, given $\bm{\tau_{<i}}$, i.e., $\mathbb{P}(\tau_{i}|\bm{\tau_{<i}})$. There are two different phases of NTP, i.e., training-time NTP via self-supervision, and inference-time NTP via autoregression.
  • Training-time NTP. During training, the model is fed the previous $\bm{\tau_{<i}}$ ground truth tokens to predict the next token, $\tau_{i}$. The model is provided with the next ground truth token $\tau_{i}$ so that it can maximize the sum of log-probabilites over the entire sequence to better predict $\tau_{i}$ for discrete actions, i.e., $\mathcal{L}_{\text{total}}=-\sum_{i=1}^{T}\log\mathbb{P}(\tau_{i}\mid\bm{\tau_{<i}})$. Notice that, unlike inference, during training the outputs generated by the models are not fed back into itself. As a result, this avoids propagating any compounding error that may result from any misprediction.
  • Inference-time NTP. During inference, NTP follows an auto-regressive process, i.e., the output generated by the model is fed back into itself to generate the next token. For example, given $\bm{\tau_{<i}}$ tokens, the model generates the next token $\tau_{i}$ that is then fed back into the model to generate the next token $\tau_{i+1}$. This process continues until a stopping criteria is met. This token generation process is inherently probabilistic, meaning that, at each step, the model outputs a probability distribution over possible next tokens $\mathbb{P}(\tau\mid\bm{\tau_{<i}})$ using either greedy selection, i.e., $\tau_{i}=\arg\max_{\tau}\mathbb{P}(\tau\mid\bm{\tau_{<i}})$, or sampling, i.e., $\tau_{i}\sim\mathbb{P}(\tau\mid\bm{\tau_{<i}})$.
Translating The NTP Paradigm into Database Systems: Many database systems problems can be framed as an NTP task. The key question is whether adopting this paradigm effectively aligns with the original objective of the task. For example, consider the Join Order Selection problem (JOS, for short) that involves 4 tables, i.e., $\mathtt{A\bowtie B\bowtie C\bowtie D}$. Framing JOS as an NTP task translates into predicting the next table to join. However, this does not align with the objective of JOS that is to determine the join order that minimizes query execution time, rather than simply determining any join order. One of the reasons behind this discrepancy is the mismatch in objectives between the domain of NLP and DBMS. DBMS optimization tasks are often goal-oriented, and are geared towards improving query performance, scalability, and resource utilization. In contrast, NLP tasks often involve generating sequences without explicitly optimizing for any particular objective. Moreover, the notion of a token in DBMS can be diverse, e.g., in the Join Order Selection, the tables can serve as the tokens, whereas in query scheduling decision problems, the queries themselves can serve as the tokens. Both of these entities are fundamentally different. Even in the Join Order Selection problem, two different database instances often have distinct sets of tables making it harder to learn meaning representation for the tables, i.e., tokens. This is unlike NLP, where the set of tokens remains fixed for a particular tokenizer. Hence, NTP does not directly translate to the field of DBMS. In the next section, we demonstrate the necessary building blocks to adopt the NTP paradigm for optimizing database systems.
3. The Building Blocks
One of the key hindrances that blocks DBMSs from benefiting from the NTP paradigm is the goal-directed nature of its optimization tasks. Besides, the notion of token in DBMSs is versatile, i.e., is context-dependent and as a whole lacks generalization capability. To this end, we need goal-conditioned sequence modeling and a redefinition of the concept of tokenization in the context of DBMSs. In this section, we discuss the building blocks in adopting the NTP paradigm for optimizing database systems.
Figure 1. The Probe and Learn (PoLe) framework.
3.1. Decision Transformer
Decision Transformer (Chen et al., 2021) (DT, for short) is a promising framework that provides the support for modeling sequences conditioned on a particular objective. DT can abstract sequence modeling as a Reinforcement Learning problem (RL, for short), where the sequential decision-making is guided by long-term rewards. It ignores traditional RL techniques, i.e., learning value functions or policies through iterative updates. Rather, it treats RL as a supervised sequence modeling, i.e., an NTP problem. Precisely, a DT is an auto-regressive sequence model that predicts the next action based on the past observations conditioned on a desired reward. By conditioning on a desired reward, DT is able to generate more effective next token predictions, contrary to the traditional NTP paradigm that operates without a specific goal. This reward-conditioned nature of DT allows it to align its next token predictions with optimal decision-making strategies. This aligns well with DBMS tasks with a pre-defined objective, e.g., minimizing query execution time.
Decision Transformers are built on top of the Transformer architecture. It represents a sequence $\bm{\tau}$ as a combination of 3 types of input tokens, i.e., reward-to-go (RTG, for short), state, and action tokens: $(\hat{r}_{0},s_{0},a_{0},\hat{r}_{1},s_{1},a_{1},\ldots,\hat{r}_{|\bm{\tau}|},s_{|\bm{\tau}|},a_{|\bm{\tau}|})$. An RTG token $\hat{r}_{t}$ denotes the future cumulative reward expected from a given timestep $t$ onward. The initial RTG $\hat{r}_{0}$ is equal to the reward of the sequence. At timestep $t$, DT uses the tokens from the latest $K$ time steps to predict the next action token $a_{t}$. $K$ is the transformer’s context length. Then, DT’s goal is to learn a policy $\pi_{\texttt{DT}}$ that estimates the probability distribution of the next action $a_{t}$, given $\bm{s_{<i}}$ and $\bm{\hat{r}_{<i}}$, i.e., $\pi_{\texttt{DT}}(a_{t}|\bm{s_{<i}},\hat{r}_{<i})$. Analogous to the standard NTP paradigm, both the state and action tokens correspond to word tokens in LLMs.
Training. The Decision Transformer (DT) is trained following the Offline RL paradigm, i.e., it learns a policy from a fixed dataset $\mathcal{D_{\texttt{OFF}}}$ without online exploration. The dataset $\mathcal{D_{\texttt{OFF}}}$ comprises a collection of experiences, i.e., state-action transitions, $(\hat{r}_{i},s_{i},a_{i})$, generated by various sources, e.g., heuristic methods, RL agents, or human demonstrations. DT does not interact with the environment, and is entirely reliant on this fixed dataset $\mathcal{D_{\texttt{OFF}}}$ to learn the best possible policy. Hence, the quality and diversity of the dataset plays a significant role in determining the performance of DT in learning an optimal policy. DT samples a mini-batch of experiences with a sequence length of $K$ from $\mathcal{D_{\texttt{OFF}}}$, and trains itself to predict the next action token using the standard L2-error (cross-entropy) for continuous (discrete) actions.
Inference. During inference, the model is fed the initial state $s_{0}$ and the desired reward, i.e., performance $\hat{r}_{0}$. Then, DT generates the next action token, $a_{0}=\pi_{\texttt{DT}}(\hat{r}_{0},s_{0})$. Once generated, $a_{0}$ executes to generate the next state token, $s_{1}\sim\mathbb{P}(s_{1}|s_{0},a_{0})$ and the next reward-to-go token, $\hat{r}_{1}=\hat{r}_{0}-R(s_{0},a_{0})$. Then, DT is fed $\hat{r}_{1},s_{1}$ to generate the next action token prediction $a_{1}$. This process repeats until all the action tokens in the sequence are generated.
Figure 2. Performance of a main-memory B+-Tree under different scheduling policies.
3.2. Tokenization
As discussed in Section 2, the discrete word tokens used in language models do not directly translate to database systems. The word tokens in language models serve as linguistic units that remain consistent across all the language tasks, and datasets. In contrast, database systems lack a universal token representation that makes it harder to adopt the NTP Paradigm.
Ideally, the universal token for database systems (DB-tokens, for short) should be able to generalize across new hardware, and workload applications. Besides, it should be computationally inexpensive to compute so that it does not affect the day-to-day database tasks. Moreover, a DB-token should be able to capture additional context, as unlike language models these action tokens lack any inherent grammar or structure. These requirements present a significant challenge in identifying an effective internal representation for databases. We propose to use as DB-tokens the hardware profiles that are generated from hardware Performance Monitoring Units (PMU, for short). Hardware profiles are computationally inexpensive to retrieve from the hardware registers, and can generalize across different hardware and workload applications, and can provide accurate hardware context that the DBMS is running on at any given time. This makes them an ideal candidate for use as DB-tokens in DBMS optimization tasks, as we demonstrate below.
What makes a hardware profile from PMUs stand out in this regard is its capability to provide a real time update of the hardware and the DBMS components at the highest possible granularity. It goes a long way in mimicking the data distribution and capturing the query workload, which have proven to be effective in designing efficient learned DBMSs, e.g., as in (Kraska et al., 2019; Pavlo et al., 2017). For example, a small region in the data space that is being queried extensively will result in a lot of cache and memory accesses in the cores that handle this data region. The number of cache accesses may further increase, if the data distribution of that region is particularly dense. Likewise, transactional queries yield less cache and memory accesses compared to the analytical queries that yield a large number of executed instructions along with higher cache and memory accesses.
Despite the ongoing evolution of the hardware landscape, the core building blocks of the hardware have remained consistent. At its core, there are compute, memory chips, interconnects and storage. Almost all modern processors provide a number of PMU counters to monitor these critical components. Thus, the PMU counters serve as an excellent fit as representative DB-tokens that capture the database system performance across varying hardware at the finest possible performance granularity. Hence, DB-tokens are an excellent choice as input tokens in an NTP paradigm for optimizing database systems.
In addition to the DB-tokens, there are action tokens to communicate policies across the DBMS and Machine Learning models. These action tokens represent modular units that make up the policy. The action tokens by themselves hold no inherent meaning. It is the DB-tokens that provide the necessary context, allowing a model to understand and learn the intricacies of DBMS optimization and how specific actions impact performance. In the next section, we describe how to connect these 2 key building blocks, i.e., Decision Transformers and the DB-tokens to develop a framework for adopting the NTP paradigm in database systems.
4. The Framework
To effectively adopt the Next Token Prediction (NTP) paradigm in database systems, we propose a novel technique termed Probe and Learn (PoLe, for short). For a given DBMS optimization task, this involves probing the hardware during query execution to generate DB-tokens. These tokens are used to train a DT to learn an optimal policy for the specified optimization task. Refer to Figure 1 for a detailed overview of the PoLe technique.
Probe: The Probe step comprises of the following 4 steps.
  • For a given optimization task with a set objective, i.e., query throughput, tail latency or scalability, we execute different policies across various hardware configurations and workloads, drawn from a diverse set of policy, hardware, and workload pools. This includes both optimal and suboptimal policies.
  • During query execution, we periodically profile the hardware to capture the behavior of important hardware components using hardware statistics, e.g., the number of instructions executed, L1-d, L1-i, Last Level Cache (LLC, for short), and branch misses.
  • Alongside the hardware profiles, we monitor the desired performance metric of the database system. These hardware profiles along with the performance metric serve as the DB-tokens. These DB-tokens are analogous to the state and RTG-tokens in DT. We also tokenize the DBMS policy to generate action tokens.
  • We periodically offload these action tokens along with the associated DB-tokens to an offline dataset for the learning phase.
Learn: The Learn step comprises of the following 2 steps.
  • We train a Decision Transformer (DT) on the collected offline dataset in a supervised manner, following the Offline Reinforcement Learning paradigm. As the DT is trained on various policies across varying hardware and workload configurations, it builds a mental model of the optimization task by gaining insights into the inherent tradeoffs and the performance improvements across different strategies, hardware and workload via the interaction of action and DB-tokens.
  • Using the trained Decision Transformer, we infer a new policy auto-regressively. The process begins by feeding DT with the initial DB-token, i.e., state and the desired objective of the specified optimization task. In turn, DT generates the next action token which is fed back to the DBMS to generate the subsequent DB-token. The DB-token is then fed back into DT to get the next action token. This iterative process continues until the complete policy is generated. Note that, during the inference process, the DT has no explicit knowledge about the hardware that the DBMS is running on, or the workload that the DBMS is executing. Hence, both the hardware, and/or the workload can be unseen. It only relies on the DB-tokens, and the inherent knowledge it has gained by being trained on a wide range of policies, hardware configurations, and workloads to generate the next action token.
Case Study: Scheduling Indexing Tasks Over NUMA Servers. To demonstrate the effectiveness of PoLe , we apply it to the main-memory index scheduling task. Modern Non Uniform Memory Access (NUMA, for short) servers are notoriously complex and heterogeneous. Each server can contain more than 1000 cores (Anandtech, 2024) spread across multiple processor dies. Data access latencies can vary significantly by up to 4$\times$ depending on the distance between the core that initiates the data request and the DIMM chip where the data resides. Additionally, concurrently running queries can cause interference by congesting on-chip and off-chip interconnects, as well as saturating hardware queues at caches and memory controllers. Thus, the performance of a DBMS index can vary significantly depending on the physical location of the core, where the queries are scheduled, and the DRAM chips where data resides.
We formulate main-memory index scheduling as an NTP task. The index is divided into into multiple index chunks, i.e., a set of index nodes. Given a set of compute cores, the goal is to predict the core on which queries should be scheduled for the next index chunk. The data placement is implicit, i.e., the data for each chunk resides in the DIMM chip associated with the assigned core. The ids of the the computing cores serve as the action tokens for scheduling. A complete policy includes a sequence of core ids. The $i$th core in the sequence represents the core assigned to process the corresponding index slice. The clock-cycles, cache, branch misses, local and remote DRAM accesses along with the query throughput constitute the DB-token.
5. Initial Evaluation
We evaluate the learned scheduling policy by PoLe against various OS policies and heuristics. The default OS policy (OS:D) is local allocation, i.e., the OS kernel places an index node on the NUMA node of the requesting CPU. The OS:I policy places the index nodes in an interleaved manner across all the NUMA nodes. For both the Shared Everything NUMA (SE:N) and the Shared Nothing (Porobic et al., 2012) (SN) strategies, nearby index slices are clustered together and placed in the same NUMA node. For OS:D, OS:I, and SE:N, the OS handles the scheduling. For SN, the scheduling strictly follows the data placement.
We run a YCSB-like benchmark (Cooper et al., 2010) on a main-memory B+-Tree on the YCSB-A workload. Initially, the index is loaded with 30M records. Each index record has a 64-bit key and a 64-bit value. Figure 2 shows the performance of the B+-Tree under different scheduling policies across different hardware. The DT in the above experiment is trained on tokens generated by executing different YCSB workloads, e.g., YCSB-A, YCSB-C, YCSB-E, etc. on a wide range of hardware. Figures 2a, 2b show that PoLe can learn scheduling policies outperforming the baselines by up to 2.78$\times$. Note that, the offline dataset that DT has been trained on contains experiences from different heuristics on the same hardware for the same workload. Thus, this shows PoLe’s ability in learning an optimal policy to improve performance. Figure 2c shows that the scheduling policy learned by PoLe can outperform the competing baselines by up to 3$\times$ on an Intel Skylake X 4 Socket 4 NUMA Node machine. Note that, the offline dataset does not contain any experience from the Intel Skylake X machine. Thus, it shows how the NTP paradigm can help generalize index performance across unseen hardware.
6. Concluding Remarks
In this paper, we lay the foundation for incorporating the Next Token Paradigm in optimizing database systems by proposing a framework that generates hardware tokens to train Decision Transformers in a data-driven framework. Whether all or which database optimization tasks can benefit from the proposed PoLe framework as a means to generalize across diverse hardware and workloads is yet to be seen. However, preliminary results on index scheduling over NUMA servers suggest that PoLe has strong potential in improving performance and adapting to diverse environments.